OpenStudy (anonymous):

how do you find the x and y intercepts of: 4x+y=5

OpenStudy (anonymous):

you have to get y alone and then you have to solve for that by = it to zero

OpenStudy (anonymous):

So for x-intercept: 4x+0=5 Divide 4 from 4x and 5? X= 5/4

OpenStudy (anonymous):

And y=5

OpenStudy (anonymous):

no hang on one minute that is not correct

OpenStudy (anonymous):

We don't have to divide 5/4.

OpenStudy (anonymous):

you have to solve for y not make y equal 0 -y=4x-5